# Required-edge Cycle Cover Problem: an ASP-Completeness Framework for Graph Problems and Puzzles

### 03 Mar 2026 v1

@@ 5.2  Choco Banana
Now we construct a reduction via the flow model as illustrated in Figure˜15. The configuration of unshaded cells with clues $37$ and $13$ is fixed by the observation above. Each block of unshaded cells containing the clue $37$ corresponds to a sink, while each block of undetermined cells corresponds to a source. These gadgets correctly realize the model, as shown below: Each sink requires unshaded cells equal to the difference between its current size and $37$, which equals its degree. On the other hand, each source supplies a total of two unshaded cells to the adjacent sinks (Figure˜16). Figure˜17 illustrates three variants of the sources, each of which restricts the admissible flow distributions as required.
@@ 5.3  Shimaguni
We construct a reduction via the flow model as illustrated in Figure˜18. Each cross-shaped region containing a clue $9$ corresponds to a sink, while each region with a clue $7$ connecting two arms of sinks corresponds to a source. These sinks and sources function as intended, as we demonstrate below: Each source supplies a total of two unshaded cells to the two connected sinks (Figure˜19). On the other hand, each sink accepts at most as many unshaded cells as its degree from its adjacent sources, due to the difference between its area and clue value. By Lemma˜11, this suffices for the reduction. Figure˜20 illustrates three variants of sources, each of which restricts the admissible flow distributions as required.
@@ 5.6  Five Cells
We construct a reduction via the flow model. Figure˜27 illustrates a unit gadget. The configuration of the surrounding yellow region is uniquely determined by the boundary conditions, specifically where the gadget abuts the board border or adjacent gadgets. The central cell of the gadget corresponds to a sink, from which four wires emanate in the four cardinal directions. Conversely, the interface between two wires of adjacent unit gadgets corresponds to a source. These gadgets correctly realize the model, as shown below: The central cell requires a total of $4$ cells from the four wires to form a pentomino, while each wire supplies a specific number of cells to the central cell (Figure˜28). Specifically, a wire connected to the board border supplies exactly $1$ cell, effectively adjusting the remaining demand of the central cell to $3$ (when abutting one border) or $2$ (when abutting two borders). For the interface between two wires, the sum of their supplies is constrained to $2$; collectively, these two connected wires function as a source.
@@
To simulate the variants of sources, we introduce two specific types of wires (Figure˜29). An incoming wire (Figure˜29(a)) can supply either $1$ or $2$ cells, whereas an exclusive wire (Figure˜29(b)) is restricted to supplying either 0 or $2$ cells. A source with exactly one incoming wire corresponds to a biased source oriented toward the incoming side. A source with two incoming wires corresponds to a fixed source, as the only pair of flow distributions summing to $2$ is $(1,1)$. A source composed of exclusive wires corresponds to an exclusive source.
@@ 6  Conclusion and Future Work
In this paper, we established the ASP-completeness of RCCP (Theorem˜6) and CCP (Corollary˜7) for planar degree-$3$ bipartite mixed graphs. This result is tight in the sense of the maximum degree, as the counting versions of these problems on max-degree-$2$ graphs are obviously polynomial-time solvable. However, the complexity for degree-$k$ graphs with $k\geq 4$ remains open. Moreover, while the input graph in Corollary˜7 is constrained such that each vertex is incident to at least one directed edge, it remains an open question whether the problem remains ASP-complete when restricted to exactly one directed edge per vertex.

### 06 Mar 2026 v2

@@ 5.2  Choco Banana
Now we construct a reduction via the flow model as illustrated in Figure˜15. The configuration of unshaded cells with clues 3737 and 1313 is fixed by the observation above. Each block of unshaded cells containing the clue 3737 corresponds to a sink, while each block of undetermined cells corresponds to a source. These gadgets correctly realize the model, as shown below: Each sink requires unshaded cells equal to the difference between its current size and 3737, which equals its degree. On the other hand, each source supplies a total of two unshaded cells to the adjacent sinks (Figure˜16). Figure˜17 illustrates three variants of the sources, each of which restricts the admissible flow distributions as required.
@@ 5.3  Shimaguni
We construct a reduction via the flow model as illustrated in Figure˜18. Each cross-shaped region containing a clue 99 corresponds to a sink, while each region with a clue 77 connecting two arms of sinks corresponds to a source. These sinks and sources function as intended, as we demonstrate below: Each source supplies a total of two unshaded cells to the two connected sinks (Figure˜19). On the other hand, each sink accepts at most as many unshaded cells as its degree from its adjacent sources, due to the difference between its area and clue value. By Lemma˜11, this suffices for the reduction. Figure˜20 illustrates three variants of sources, each of which restricts the admissible flow distributions as required.
@@ 5.6  Five Cells
We construct a reduction via the flow model. Figure˜27 illustrates a unit gadget. The configuration of the surrounding yellow region is uniquely determined by the boundary conditions, specifically where the gadget abuts the board border or adjacent gadgets. The central cell of the gadget corresponds to a sink, from which four wires emanate in the four cardinal directions. Conversely, the interface between two wires of adjacent unit gadgets corresponds to a source. These gadgets correctly realize the model, as shown below: The central cell requires a total of 44 cells from the four wires to form a pentomino, while each wire supplies a specific number of cells to the central cell (Figure˜28). Specifically, a wire connected to the board border supplies exactly 11 cell, effectively adjusting the remaining demand of the central cell to 33 (when abutting one border) or 22 (when abutting two borders). For the interface between two wires, the sum of their supplies is constrained to 22; collectively, these two connected wires function as a source.
@@
To simulate the variants of sources, we introduce two specific types of wires (Figure˜29). An incoming wire (Figure˜29(a)) can supply either 11 or 22 cells, whereas an exclusive wire (Figure˜29(b)) is restricted to supplying either 0 or 22 cells. A source with exactly one incoming wire corresponds to a biased source oriented toward the incoming side. A source with two incoming wires corresponds to a fixed source, as the only pair of flow distributions summing to 22 is $(1,1)$. A source composed of exclusive wires corresponds to an exclusive source.
@@ 6  Conclusion and Future Work
In this paper, we established the ASP-completeness of RCCP (Theorem˜6) and CCP (Corollary˜7) for planar degree-33 bipartite mixed graphs. This result is tight in the sense of the maximum degree, as the counting versions of these problems on max-degree-22 graphs are obviously polynomial-time solvable. However, the complexity for degree-$k$ graphs with $k\geq 4$ remains open. Moreover, while the input graph in Corollary˜7 is constrained such that each vertex is incident to at least one directed edge, it remains an open question whether the problem remains ASP-complete when restricted to exactly one directed edge per vertex.